# Optimal Oblivious Load-Balancing for Sparse Traffic in Large-Scale Satellite Networks

### v3

@@ Appendix G Proof of Lemma 4
In the minimal subset $S^{*}=S_{r}(0)$, observe that each node in $S_{r}(0)$ contributes to at least 22 edges into the cut set. This tells us that there are at least two edge-disjoint paths emanating from each node in $S_{r}(0)$. A similar argument can be made for edge disjoint paths terminating at $S_{r}(t)$ by considering $S^{*}$ to be $S_{r}(t)$. ∎
@@ Appendix I Proof of Worst-Case Lower Bound in General $N\times M$ Tori
Next, we define $B^{(\lambda_{1},\lambda_{2})}_{r}(i)$ to be the ball around node $i$ containing nodes which are under $r$ “generalized” hops away from ii. In other words,
@@
We consider the dual of the problem (18). We focus on the case where NN is even for simplicity. The case when NN is odd is similar, with minor modifications. From weak duality, problem (18) can be lower bounded by
@@
In this case, the no overlap condition is again equivalent to $r_{2}<t_{x}$ and $r_{1}<t_{y}$. The MinCut condition tells us that the minimal cut between the stems is strictly lesser than $2r_{1}+2r_{2}$, which is the number of required edge-disjoint paths from the source stem to the destination stem in the previous case. Additionally, when $\textsc{MinCut}(S_{r_{1},r_{2}}(0),S_{r_{1},r_{2}}(t))<4r_{1}+4r_{2}$, the MinCut is in fact close to the bisection cut, $\min(2N,2M)$. Simply put, this happens when the $N\times M$ network is skewed along one dimension, i.e., either $M$ or NN is much greater than the other, and so the smallest cut set separating the stems $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ is smaller than $4r_{1}+4r_{2}$. Without loss of generality, we assume $N<M$, i.e., the torus is longer in the horizontal direction $e_{2}$ and shorter in the vertical direction. In the case when $N>M$, the same description of the routing policy is the same, however, with the roles of the horizontal and vertical links reversed. We describe the phases of the GLLB when $N<M$ as follows.
@@
In this case, the overlap condition is equivalent to $r_{2}\geq t_{x}$ or $r_{1}\geq t_{y}$. Without loss in generality, there can only be the two types of overlap between stems, Type-I and type-II, similar to Fig. 10 and Fig. 11. This is similar to the case of overlap described Appendix F, with the only difference being in the distribution of traffic among the true stems.
J-C1 Type-I Overlap
In Type-I overlap, the stems $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ overlap only along one of the four legs of the stem. Under Type-I overlap, we define modified stems $S^{I}_{r_{1},r_{2}}(0)$ and $S^{I}_{r_{1},r_{2}}(t)$,
$S^{I}_{r_{1},r_{2}}(0)$ $=\{j\in V|i\in S_{r_{1},r_{2}}(0)\text{ and }\Delta(0,j)\leq\Delta(t,j)\}$
$S^{I}_{r_{1},r_{2}}(t)$ $=\{j\in V|i\in S_{r_{1},r_{2}}(t)\text{ and }\Delta(t,j)\leq\Delta(0,j)\}.$
J-C2 Type-II Overlap
In Type-II overlap, the stems $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ overlap exactly at two nodes at two of the four legs of the stem respectively. Under Type-II overlap, let $(0,h_{x})$ and $(h_{y},0)$ be the indices of the nodes that lie in both $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$. We define modified stems $S^{II}_{r_{1},r_{2}}(0)$ and $S^{II}_{r_{1},r_{2}}(t)$,
$S^{II}_{r_{1},r_{2}}(0)$ $=\{j\in V|\;j\in S_{r_{1},r_{2}}(0)\text{ and }j_{x}\leq h_{x},\;\;j_{y}\leq h_{y}\}$
$S^{II}_{r_{1},r_{2}}(t)$ $=\{j\in V|\;j\in S_{r_{1},r_{2}}(t)\text{ and }j_{x}\leq t_{x}-h_{x},\;\;j_{y}\leq t_{y}-h_{y}\}.$
Once we define the modified stem for each type of overlap, the modifications in each routing phase is fairly straightforward. Henceforth, $S^{\prime}_{r_{1},r_{2}}(0)$ and $S^{\prime}_{r_{1},r_{2}}(t)$ refer to the modified stems in each case of overlap.
$\bullet$ Modification in the first phase: Distribute $1/2(r_{1}+r_{2})$ fraction of traffic to each of the nodes in the stems $S^{\prime}_{r_{1},r_{2}}(0)$, except the last node in each of the legs. If the node is the last node in its leg and is $h$ hops away, it receives $(r_{\ell}-h+1)/2(r_{1}+r_{2})$ fraction of the traffic, depending on which leg has the overlap. Here, $\ell=1$ for a vertical leg and $\ell=2$ for a horizontal leg.
$\bullet$ Modification in the second phase: Route traffic from each node in the stem $S^{\prime}_{r_{1},r_{2}}(0)$ to each node in the stem $S^{\prime}_{r_{1},r_{2}}(t)$ (or $S^{\prime}_{r_{1},r_{2}}(t)$ along edge-disjoint paths, such that each node in $S^{\prime}_{r_{1},r_{2}}(0)$ sends equal traffic along 2 unique edge-disjoint paths and each node in $S^{\prime}_{r_{1},r_{2}}(t)$ receives traffic along 2 edge-disjoint paths. Each path carries $1/4(r_{1}+r_{2})$ fraction of traffic. If there is residual traffic at a node in $S^{\prime}_{r_{1},r_{2}}(0)$, then there must exist a node neighboring it that is in $S^{\prime}_{r_{1},r_{2}}(t)$. The residual traffic must be directly sent to that neighbor. If a node belongs to both $S^{\prime}_{r_{1},r_{2}}(0)$ and $S^{\prime}_{r_{1},r_{2}}(t)$, then the node needs to do nothing in the second phase.
$\bullet$ Modification in the third phase: Aggregate traffic from the nodes in stem $S^{\prime}_{r_{1},r_{2}}(t)$ to the final destination of the traffic, $t$, along the shortest path. As a consequence of the first two stages, $1/2(r_{1}+r_{2})$ fraction of traffic needs to be aggregated from each node in $S^{\prime}_{r_{1},r_{2}}(t)$ except the last nodes in each leg. A fraction of $(r_{\ell}-h+1)/4r$ of the traffic will be aggregated from the last node in each leg, where the last node is $h$ hops away. Here, $\ell=1$ for a vertical leg and $\ell=2$ for a horizontal leg.
@@
In this case, the overlap condition is equivalent to $r_{2}\geq t_{x}$ or $r_{1}\geq t_{y}$. As in the previous case, there are only two possible types of overlap, Type-I and Type-II, which are described in the previous section. Once we define the modified stem for each type of overlap, the modifications in each routing phase is fairly straightforward. Henceforth, $S^{\prime}_{r_{1},r_{2}}(0)$ and $S^{\prime}_{r_{1},r_{2}}(t)$ refer to the modified stems in each case of overlap.
$\bullet$ Modification in the first phase: We distribute $1/N$ fraction of traffic to each of the nodes in vertical legs of the stem $S^{\prime}_{r_{1},r_{2}}(0)$, except the last node in each of the legs. We distribute $(1/r_{2}-2r_{1}/r_{2}N)$ fraction of traffic to each of the nodes in horizontal legs of the stem $S^{\prime}_{r_{1},r_{2}}(0)$, except the last node in each of the legs. If the node is the last node in a vertical leg and is $h$ hops away, it receives $(r_{1}-h+1)/N$ fraction of the traffic. If the node is the last node in a horizontal leg and is $h$ hops away, it receives $(r_{2}-h+1)(1/r_{2}-2r_{1}/r_{2}N)$ fraction of the traffic.
$\bullet$ Modification in the second phase: In this phase, we route traffic from the source stem $S^{\prime}_{r_{1},r_{2}}(0)$ to the destination stem $S^{\prime}_{r_{1},r_{2}}(t)$ such that the vertical and horizontal links outside the stem carry a maximum fraction of $\lambda_{1}$ and $\lambda_{2}$ of the traffic respectively. We choose $\lambda_{2}=1/2N$ and $\lambda_{1}=(1/2r_{2}-r_{1}/r_{2}N)$. The intuition for this is the same as in Case 2 in section J-B. The traffic from $\lceil\lambda_{2}/\lambda_{1}\rceil$ vertical edges combine at some intermediate nodes in order to maintain the maximum of $\lambda_{2}$ fraction of traffic along horizontal edges. It is always possible to find such a set of paths based on the choice of our $\lambda_{1}$ and $\lambda_{2}$, and, due to the geometry of the network. If there is residual traffic at a node in $S^{\prime}_{r_{1},r_{2}}(0)$, then there must exist a node neighboring it that is in $S^{\prime}_{r_{1},r_{2}}(t)$. The residual traffic must be directly sent to that neighbor. If a node belongs to both $S^{\prime}_{r_{1},r_{2}}(0)$ and $S^{\prime}_{r_{1},r_{2}}(t)$, then the node needs to do nothing in the second phase.
$\bullet$ Modification in the third phase: Aggregate traffic from the nodes in stem $S^{\prime}_{r_{1},r_{2}}(t)$ to the final destination of the traffic, $t$, along the shortest path. As a consequence of the first two stages, $1/N$ and $(1/r_{2}+r_{1}/r_{2}N)$ fraction of traffic needs to be aggregated from each node in the vertical and horizontal legs of $S^{\prime}_{r_{1},r_{2}}(t)$ respectively, excepting the last nodes in each leg. If the node is the last node in a vertical leg and is $h$ hops away, $(r_{1}-h+1)/N$ fraction of the traffic needs to be aggregated from it. If the node is the last node in a horizontal leg and is $h$ hops away, $(r_{2}-h+1)(1/r_{2}-2r_{1}/r_{2}N)$ fraction of the traffic needs to be aggregated from it.
@@
Without loss in generality, assume $N>M$. We choose $r_{1}=\left\lceil\sqrt{\frac{c_{1}k}{2c_{2}}}\right\rceil$ and $r_{2}=\left\lceil\sqrt{\frac{c_{2}k}{2c_{1}}}\right\rceil$. Since $k\leq\min\left(c_{2}N^{2}/c_{1},c_{1}M^{2}/c_{2}\right)$, there will be $2r_{1}+2r_{2}$ nodes in each of the stems. Let us define
@@
Based on this routing policy, the load on any vertical link is at most $N/4c_{1}$. This is because the load due uniformly distributing unit traffic from each node in a ring of size NN with link capacity $c_{1}$ is $N/8c_{1}$ [1, 4]. This load appears twice, once due to distribution phase and once due to the aggregation phase.

### v4

@@ Appendix G Proof of Lemma 4
In the minimal subset $S^{*}=S_{r}(0)$, observe that each node in $S_{r}(0)$ contributes to at least $2$ edges into the cut set. This tells us that there are at least two edge-disjoint paths emanating from each node in $S_{r}(0)$. A similar argument can be made for edge disjoint paths terminating at $S_{r}(t)$ by considering $S^{*}$ to be $S_{r}(t)$. ∎
@@ Appendix I Proof of Worst-Case Lower Bound in General $N\times M$ Tori
Next, we define $B^{(\lambda_{1},\lambda_{2})}_{r}(i)$ to be the ball around node $i$ containing nodes which are under $r$ “generalized” hops away from $i$. In other words,
@@
We consider the dual of the problem (18). We focus on the case where $N$ is even for simplicity. The case when $N$ is odd is similar, with minor modifications. From weak duality, problem (18) can be lower bounded by
@@
In this case, the no overlap condition is again equivalent to $r_{2}<t_{x}$ and $r_{1}<t_{y}$. The MinCut condition tells us that the minimal cut between the stems is strictly lesser than $2r_{1}+2r_{2}$, which is the number of required edge-disjoint paths from the source stem to the destination stem in the previous case. Additionally, when $\textsc{MinCut}(S_{r_{1},r_{2}}(0),S_{r_{1},r_{2}}(t))<4r_{1}+4r_{2}$, the MinCut is in fact close to the bisection cut, $\min(2N,2M)$. Simply put, this happens when the $N\times M$ network is skewed along one dimension, i.e., either $M$ or $N$ is much greater than the other, and so the smallest cut set separating the stems $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$ is smaller than $4r_{1}+4r_{2}$. Without loss of generality, we assume $N<M$, i.e., the torus is longer in the horizontal direction $e_{2}$ and shorter in the vertical direction. In the case when $N>M$, the same description of the routing policy is the same, however, with the roles of the horizontal and vertical links reversed. We describe the phases of the GLLB when $N<M$ as follows.
@@
In this case, the overlap condition is equivalent to $r_{2}\geq t_{x}$ or $r_{1}\geq t_{y}$. This is similar to the case of overlap described Appendix F. In this case, there exist nodes that lie in both $S_{r_{1},r_{2}}(0)$ and $S_{r_{1},r_{2}}(t)$. The modifications in each routing phase is fairly straightforward.
$\bullet$ First phase: Distribute $1/2(r_{1}+r_{2})$ fraction of traffic to each of the nodes in the stems $S_{r_{1},r_{2}}(0)$ along shortest paths.
$\bullet$ Second phase: Route traffic from each node in the stem $S_{r_{1},r_{2}}(0)$ to each node in the stem $S_{r_{1},r_{2}}(t)$ along edge-disjoint paths, such that each node in $S_{r_{1},r_{2}}(0)$ sends equal traffic along 2 unique edge-disjoint paths and each node in $S_{r_{1},r_{2}}(t)$ receives traffic along 2 edge-disjoint paths. If the node in $S_{r_{1},r_{2}}(0)$ also lies in $S_{r_{1},r_{2}}(t)$, no traffic needs to be explicitly forwarded, as we have already reached the destination stem. Additionally, the paths in the second phase must not use the edges used in the first and third phases. Each path carries $1/4(r_{1}+r_{2})$ fraction of traffic.
$\bullet$ Third phase: Aggregate traffic from the nodes in stem $S_{r_{1},r_{2}}(t)$ to the final destination of the traffic, $t$, along the shortest path. As a consequence of the first two stages, $1/2(r_{1}+r_{2})$ fraction of traffic needs to be aggregated from each node in $S_{r_{1},r_{2}}(t)$.
Due to overlaps in the stem sets, there could be loopy paths that could formed. There could also be edges used in the first and third phase that are used twice (once in the first phase and once in the third phase). As a final step, the final flow in the forward direction needs to be updated to remove loops or redundant use of edges
@@
In this case, the overlap condition is equivalent to $r_{2}\geq t_{x}$ or $r_{1}\geq t_{y}$. The modifications are similar to the previous case, and are fairly straightforward.
$\bullet$ First phase: We distribute $1/N$ fraction of traffic to each of the nodes in vertical legs of the stem $S_{r_{1},r_{2}}(0)$, except the last node in each of the legs. We distribute $(1/r_{2}-2r_{1}/r_{2}N)$ fraction of traffic to each of the nodes in horizontal legs of the stem $S_{r_{1},r_{2}}(0)$.
$\bullet$ Second phase: In this phase, we route traffic from the source stem $S_{r_{1},r_{2}}(0)$ to the destination stem $S_{r_{1},r_{2}}(t)$ such that the vertical and horizontal links outside the stem carry a maximum fraction of $\lambda_{1}$ and $\lambda_{2}$ of the traffic respectively. We choose $\lambda_{2}=1/2N$ and $\lambda_{1}=(1/2r_{2}-r_{1}/r_{2}N)$. The intuition for this is the same as in Case 2 in section J-B.
If the node in $S_{r_{1},r_{2}}(0)$ also lies in $S_{r_{1},r_{2}}(t)$, no traffic needs to be explicitly forwarded, as we have already reached the destination stem. Additionally, the paths in the second phase must not use the edges used in the first and third phases.
The traffic from $\lceil\lambda_{2}/\lambda_{1}\rceil$ vertical edges combine at some intermediate nodes in order to maintain the maximum of $\lambda_{2}$ fraction of traffic along horizontal edges. It is always possible to find such a set of paths based on the choice of our $\lambda_{1}$ and $\lambda_{2}$, and, due to the geometry of the network.
$\bullet$ Third phase: Aggregate traffic from the nodes in stem $S_{r_{1},r_{2}}(t)$ to the final destination of the traffic, $t$, along the shortest path. As a consequence of the first two stages, $1/N$ and $(1/r_{2}+r_{1}/r_{2}N)$ fraction of traffic needs to be aggregated from each node in the vertical and horizontal legs of $S_{r_{1},r_{2}}(t)$ respectively, excepting the last nodes in each leg.
Due to overlaps in the stem sets, there could be loopy paths that could formed. There could also be edges used in the first and third phase that are used twice (once in the first phase and once in the third phase). As a final step, the final flow in the forward direction needs to be updated to remove loops or redundant use of edges
@@
Without loss in generality, assume $M>N$. We choose $r_{1}=\left\lceil\sqrt{\frac{c_{1}k}{2c_{2}}}\right\rceil$ and $r_{2}=\left\lceil\sqrt{\frac{c_{2}k}{2c_{1}}}\right\rceil$. Since $k\leq\min\left(c_{2}N^{2}/c_{1},c_{1}M^{2}/c_{2}\right)$, there will be $2r_{1}+2r_{2}$ nodes in each of the stems. Let us define
@@
Based on this routing policy, the load on any vertical link is at most $N/4c_{1}$. This is because the load due uniformly distributing unit traffic from each node in a ring of size $N$ with link capacity $c_{1}$ is $N/8c_{1}$ [1, 4]. This load appears twice, once due to distribution phase and once due to the aggregation phase.